Given that tan x°=3/7 for which x is an acute angle, find without using using mathematical tables or a calculator, cos(90-x)°
Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Given that tan x°=3/7 for which x is an acute angle, find without using using mathematical tables or a calculator, cos(90-x)°
-------------
Make right triangle with sides of y = 3 and x = 7
r = sqrt(3^2 + 7^2) = sqrt(58)
cos(90-x) = sin(x) = y/r = 3/sqrt(58)

Answer by ikleyn(52787)

.

Notice that without using tables or a calculator,  we can find an exact  EXPRESSION  for  cos(90°-x),  ONLY,
but not its value itself.


I am telling it to explain you that the formulation of the problem is incorrect and should be changed.